# LAMBench: A Benchmark for Large Atomic Models

### v1

@@ IV Method
IV.1 Models
To evaluate the thoroughness and discriminative power of the LAMBench benchmark, we test a series of LAMs, as listed in Table 4. Most models follow a single-task, single-fidelity training strategy, using datasets curated under consistent DFT settings—primarily from the domain of Inorganic Materials. Exceptions include SevenNet-MF-OMPA and DPA-2.4-7M. SevenNet-MF-OMPA employs multi-fidelity training on the OMat24[50], MPtrj, and sAlex[50] datasets, achieving high accuracy despite heterogeneity in DFT settings, though still focused on inorganic systems. In contrast, DPA-2.4-7M adopts a multitask training approach using the OpenLAM dataset collection (Supporting Information LABEL:tab:si-datasets), which spans a broad range of chemical and material systems—including catalysis, small molecules, reactions, and biomolecules—under diverse DFT settings. Detailed hyperparameter information for the models can be found in Supporting Information Table S-1. The multitask training of DPA-2.4-7M produces 31 prediction heads, each corresponding to a distinct training task within OpenLAM. Unless otherwise noted, the prediction head trained by the MPtrj dataset is used by default.
@@ IV.1 Models
GRACE-2L-OAM[62] 15.3M MPtrj No 6.0
IV.2 LAMBench Implementation
@@ IV.3 Generalizability metrics
[…]
$\hat{M}^{m}_{k,p,i}=\frac{M^{m}_{k,p,i}}{M^{\mathrm{dummy}}_{k,p,i}}$ (1)
This baseline model predicts energy based solely on the chemical formula, disregarding any structural details, thereby providing a reference point for evaluating the improvement offered by more sophisticated models.
For each domain, we compute the log-average of normalized metrics across all datasets within this domain by
@@
For the force field generalizability tasks, we adopt RMSE as the primary error metric. The prediction types include energy and force, with weights assigned as $w_{E}=w_{F}=0.5$. When periodic boundary conditions are assumed and virial labels are available, virial predictions are also considered. In this scenario, the prediction weights are adjusted to $w_{E}=w_{F}=0.45$ and $w_{V}=0.1$. The resulting error metric is referred to as $\bar{M}^{m}_{\mathrm{FF}}$.
For the domain-specific property calculation tasks, we adopt the MAE as the primary error metric. In the Inorganic Materials domain, the MDR phonon benchmark predicts maximum phonon frequency, entropy, free energy, and heat capacity at constant volume, with each prediction type assigned a weight of 0.25. In the Small Moleculesdomain, the TorsionNet 500 benchmark predicts the torsion profile energy, torsional barrier height, and the number of molecules for which the model’s prediction of the torsional barrier height has an error exceeding 1 kcal/mol. Each prediction type in this domain is assigned a weight of $\frac{1}{3}$. The resulting error metric is denoted as $\bar{M}^{m}_{\mathrm{PC}}$.
IV.4 Applicability metrics
The applicability metrics incorporate both efficiency and stability and are computed differently from the generalizability metrics due to the absence of a dummy model baseline. To evaluate efficiency, we define an efficiency metric, $M_{E}^{m}$ , by normalizing the average inference time (with unit $\mathrm{\mu s/atom}$), $\bar{\eta}^{m}$, of a given LAM measured over 1600 configurations with respect to an artificial reference value, thereby rescaling it to a range between zero and positive infinity. A larger value indicates the higher efficiency.
$M_{\mathrm{E}}^{m}=\frac{\eta^{0}}{\bar{\eta}^{m}},\quad\eta^{0}=100\ \mathrm{\mu s/atom},\quad\bar{\eta}^{m}=\frac{1}{1600}\sum_{i}^{1600}\eta_{i}^{m},$ (5)
@@ IV.4 Applicability metrics
This result is bounded within the range $[0,+\infty)$, where a lower value signifies greater stability.
@@ Acknowledgments
Module Hyperparameter DPA-2.4-7M Repinit nsel 120 neuron 3$\times$64 three_body_neuron 3$\times$32 Repformer nsel 40 nlayers 6 g1_dim 384 g2_dim 96 Activation tanh Training batch size “auto:256” nGPUs 120 $\times$ A800-40GB start lr 1e-3 stop lr 1e-5 precision float32 steps trained 8M
@@
Table S-2: Summary of datasets including dataset name, frame numbers, data cleaning, training average atoms, and DFT level.
@@
Dataset Name Frames Train Avg Atoms Data Cleaning Level of Theory
@@
Alex2D[65] trn: 1,223,831 val: 135,810 9.9 Exclude frames with energies $>$ 0 eV/atom, maximum absolute force $>$ 5 eV/Å, maximum absolute virial/atom $>$ 8 eV/atom; Trajectory first/last 3 frames + every 10 frame PBE/PAW, 520 eV, 0.4 Å-1
@@
MPTraj[13] trn: 1,401,956 val: 110,918 31.35 According to the distribution difference between MACE-MP-0 prediction and label, the union set of top 0.1% with the largest uncorrected energy difference and the top 0.1% with the largest force difference were excluded, resulting in a total of 3123 frames excluded. PBE (+U)/PAW, 520 eV, 0.04 Å-1
OC20M[66] trn: 20,000,000 val: 999,866 73.1 - rPBE/PAW, 350eV
@@
OC22[67] trn: 8,194,770 val: 394,727 79.8 Exclude frames with energies $>$ 0eV/atom, maximum absolute force $>$ 10eV/Å PBE (+U)/PAW, 350 eV, 0.157 Å-1
@@
ODAC23[68] trn: 2,682,332 val: 63,623 203 Exclude frames with energy/atom $>$ 0.5 eV/atom or $<$ -0.2 eV/atom, and maximum absolute force $>$ 25 eV/Å; Trajectory first/last 3 frames + every 20th frame PBE-D3/PAW, 600 eV, gamma
@@
OMAT24[50] trn: 100,568,820 val: 1,074,647 18.6 Exclude frames with energies $>$ 0 eV/atom, $<$ -25eV/atom, maximum absolute force $>$ 50eV/Å PBE (+U)/PAW
SPICE2[69] trn: 1,621,168 val: 180,089 35.6 Exclude frames with energies $<$ -10000 eV/atom, maximum absolute force $>$ 15eV/Å $\omega$B97M-D3(BJ)/def2-TZVPPD
Transition1x[70] trn: 7,632,328 val: 967,454 13.91 - $\omega$B97X/6-31G(d)
UniPero[71] trn: 14,487 val: 1,357 53.64 - PBEsol/LCAO, 1360 eV, 0.189 Å-1
Dai2023Alloy[72] trn: 71,482 val: 1,240 20.99 - PBE/Norm-conserving, 1360 eV, 0.094 Å-1
Zhang2023Cathode[73] trn: 88,692 val: 9,695 46.28 - PBE (+U)/PAW, 520 eV, 0.25 Å-1
Gong2023Cluster[74] trn: 143,418 val: 15,331 23.67 - PBE-D3/TZV2P, 400-1000 Ry
Li2025APEX[75] trn: 24,097 val: 100 24.03 - PBE/NC, 1360 eV, 0.15 Å-1
Shi2024Electrolyte[76] trn: 65,393 val: 3,438 192.69 - PBE-D3, 800 Ry
Shi2024SSE[77] trn: 125,083 val: 6,587 48.33 - PBE-sol/LCAO, 1360 eV, 0.28 Å-1
Yang2023ab[78] trn: 1,379,956 val: 24,257 33.47 - $\omega$B97X-D/6-31G**
Li2025General[79] trn: 14,024,587 val: 1,558,260 17.8 Exclude maximum absolute force $>$ 20eV/Å GFN2-xTB
Huang2021Deep-PBE[80] trn: 17,582 val: 886 218.59 - PBE/PAW, 650 eV, 0.26 Å-1
Liu2024Machine[81] trn: 215,481 val: 23,343 70.36 - PBE/LCAO, 1360 eV, 0.151 Å-1
Zhang2021Phase[82] trn: 46,077 val: 2,342 172.3 - SCAN/PAW, 1500 eV, 0.5 Å-1
Jiang2021Accurate[83] trn: 138,194 val: 3,965 23.16 - PBE/PAW, 650 eV, 0.1 Å-1
Chen2023Modeling[84] trn: 6,449 val: 276 16.94 - SCAN/PAW, 650 eV, 0.1 Å-1
Unke2019PhysNet[85] trn: 2,594,609 val: 136,571 21.38 - revPBE-D3(BJ)/def2-TZVP
Wen2021Specialising[86] trn: 10,054 val: 474 20.02 - PBE/PAW, 650 eV, 0.1 Å-1
Wang2022Classical[87] trn: 14,935 val: 738 24.47 - PBE/PAW, 650 eV, 0.1 Å-1
Wang2022Tungsten[88] trn: 42,297 val: 2,100 24.91 - PBE/PAW, 600 eV, 0.16 Å-1
Wu2021Deep[89] trn: 27,660 val: 917 96 - PBE/PAW, 600 eV, 2$\times$2$\times$2 kpt
Huang2021Deep-PBEsol[80] trn: 7,502 val: 384 160.8 - PBE-sol/PAW, 650 eV, 0.26 Å-1
Wang2021Generalizable[90] trn: 64,239 val: 2,256 21.85 - PBE-D3/PAW, 650 eV, 0.1 Å-1
Wu2021Accurate[91] trn: 11,621 val: 568 21.73 - PBE/PAW, 700 eV
Tuo2023Hybrid[92] trn: 48,078 val: 2,530 45.11 - PBE-D3(BJ)/PAW, 500 eV, 0.16 Å-1
@@
Table S-3: Phonon related property prediction on MDR phonon benchmark.
@@
Table S-4: Torsional MAE and MAEB errors between LAM predictions and its reference DFT labels on the TorsionNet-500 Benchmark.
@@
DPA-2.4-7M 0.727 1.220 267
MACE-MP-0 1.653 2.425 354
MACE-MPA-0 1.486 2.179 339
Orb-v2 1.238 2.067 345
SevenNet-l3i5 1.113 1.641 300
SevenNet-MF-ompa 1.599 2.309 354
MatterSim-v1-5M 1.348 2.240 345
GRACE-2L-OAM 1.387 1.992 303
dummy 2.501 5.877 494
@@
S-3 Downsampling Details for OOD Datasets
We applied a variety of down-sampling strategies across eleven benchmark datasets to reduce computational cost while preserving representativeness.
The datasets employed in this study underwent various downsampling procedures to ensure manageable sizes while maintaining representative distributions. For the ANI-1x dataset, 10% of frames were randomly sampled from the validation set using a fixed random seed 42. The Lopanitsyna2023Modeling_B dataset was used in its entirety without downsampling, while its surface counterpart Lopanitsyna2023Modeling_A utilized the validation split provided by the original authors. The combined Dai2024Deep dataset was reduced to 5% of its original size through random sampling with seed 0. For the MD22 dataset, 5% of frames were randomly sampled from systems of short peptides (Ac-Ala3-NHMe), docosahexaenoic acid, stachyose, and DNA base pairs (AT-AT and AT-AT-CG-CG), with random seed set to 0 and units fixed. The Gasteiger2020Fast dataset retained its original test split without any downsampling. The Guan2022Benchmark dataset was 5% subset randomly sampled (seed 0) from the pre-cleaned data which excluding frames with energy $>-0.5~{}\mathrm{eV/atom}$ and maximum force $>30~{}\mathrm{eV/\text{\AA}}$. The Zhang2019Bridging dataset excluded frames with energy $<0.53~{}\mathrm{eV/atom}$ and maximum forces $>5~{}\mathrm{eV/\text{\AA}}$. The Batzner2022equivariant dataset was downsampled to 5% (seed 0) after removing frames with energy $<-50~{}\mathrm{eV/atom}$ and maximum forces $>10~{}\mathrm{eV/\text{\AA}}$. The AIMD-Chig dataset underwent downsampling at 0.5% (seed 0). The Zhang2024Active dataset was randomly downsampled to 5% (seed 0) after removing frames with energy $>-2.5~{}\mathrm{eV/atom}$ and maximum forces $>25~{}\mathrm{eV/\text{\AA}}$. The Villanueva2024Water dataset was used in full after removing frames with energy ($<-6~{}\mathrm{eV/atom}$) and force ($<2~{}\mathrm{eV/\text{\AA}}$). The Subalex_9k dataset was created by randomly sampling 9157 frames (seed 42) from pre-cleaned data that excluded frames with energy $<-25~{}\mathrm{eV/atom}$ and maximum forces $>3~{}\mathrm{eV/\text{\AA}}$ from sAlex. The WBM_25k dataset was downsampled to 25696 frames (seed 42). The Sours2023Applications dataset was reduced to 5% (seed 0) after removing frames with energy $>-5~{}\mathrm{eV/atom}$. The Torres2019Analysis (Ca_Battery) was used in all GGA-labeled dataset with no virials. The Vandermause2022Active was used in its full set of DFT labels.
S-4 Structures for Conservativeness Benchmark
@@
S-5 Inference Efficiency Convergence Test
@@
S-6 The Coordination Number Analysis of Configurations in Efficiency Benchmarks
@@
S-7 Property Fine-tuning Batch Size Effect
As mentioned earlier, prediction accuracy is strongly correlated with the training batch size due to the limited size of the fine-tuning dataset. The default workflow utilizes NVIDIA L4 GPUs for cost management purposes, which results in a notable decline in performance. However, the relative performance trends among models remain consistent. For the MP $E_{\text{form}}$ and MP Gap tasks, the reported accuracy may not reflect the fully converged results due to insufficient training epochs.
Table S-5: Comparison of property fine-tuning accuracy on NVIDIA L4-24GB and A800-40GB GPU. For MP $E_{\text{form}}$ and MP Gap, the batch size on the L4 was set to a quarter of that used on the A100, while for all other tasks, it was set to half.
Tasks (Units) DPA-2.4-7M L4 Finetune DPA-2.4-7M A800 Finetune Estimated Epochs11footnotemark: 1 L4 |||| A800 MP $E_{\text{form}}$ (meV/atom) 31.1 24.1 80 |||| 320 MP Gap (eV) 0.285 0.206 100 |||| 400 JDFT2D (meV/atom) 46.93 34.09 28500 |||| 57000 Phonons (cm-1) 41.10 34.00 14000 |||| 28000 Dielectric (unitless) 0.431 0.300 3000 |||| 6000 Log KVRH ($\log_{10}\text{GPA}$) 0.062 0.052 1300 |||| 2600 Log GVRH ($\log_{10}\text{GPA}$) 0.079 0.068 1300 |||| 2600 Perovskites (eV/unitcell) 0.061 0.032 1350 |||| 2700
11footnotemark: 1
The DPA models employ a dynamic batch size to optimize training efficiency, making it difficult to precisely determine the number of training epochs.
S-8 Stability Test Results
Table S-6: NVE molecular dynamics simulations over nine atomic systems.

### v2

@@ IV Method
IV.1 computational details
Energy and force labels of ANI-1x, MD22 and AIMD-Chig sets are calculated using density functional theory (DFT) at the level of generalized gradient approximation (GGA) in the form of Perdew, Burke and Ernzerhof (PBE) functional[69][70] with 6-31G(d) basis sets in Gaussian 16 Rev. C.01[71]. self-consistent field (SCF) convergence requires the root mean squared (RMS) change in the density matrix smaller than $1\times 10^{-8}$, meanwhile the maximum change in the density matrix smaller than $1\times 10^{-6}$. Integration grids are set to 99 radial shells per atom and 590 angular points per shell. Calculations for ANI-1x and MD22 systems adopt the neutral and closed-shell singlet states, which is consistent with the original datasets. Calculations for AIMD-Chig systems adopt the -2 charged and closed-shell singlet states, given all molecules in this set are chignolin dianions.
The energies, interatomic forces and virials labels of the Lopanitsyna2023Modeling and Mazitov2024Surface data on were obtained from density functional theory (DFT) calculations performed using the Vienna Ab initio Simulation Package (VASP). The electronic structure calculations employed the generalized gradient approximation with the PBE functional for exchange-correlation effects, adopting parameter settings consistent with the Materials Project database, including a plane-wave cutoff energy of 520 eV. The electronic self-consistency cycle was converged to $5\times 10^{-5}$ eV. For Brillouin zone integration, we used the Methfessel-Paxton smearing method (ISMEAR = 1) with a smearing width of 0.05 eV, which is particularly suitable for metallic systems like high-entropy alloys. The k-point mesh was generated automatically based on a reciprocal density of 64, corresponding to a k-spacing of approximately 0.5 $^{-1}$, ensuring sufficient sampling accuracy while maintaining computational efficiency. Core-electron interactions were treated using projector-augmented wave pseudopotentials, and all calculations were performed without spin polarization to maintain consistency across the diverse compositional space of our high-entropy alloy systems.
IV.2 Models
To evaluate the thoroughness and discriminative power of the LAMBench benchmark, we test a series of LAMs, as listed in Table 4. Most models follow a single-task, single-fidelity training strategy, using datasets curated under consistent DFT settings—primarily from the domain of Inorganic Materials. Exceptions include SevenNet-MF-OMPA and DPA models. SevenNet-MF-OMPA employs multi-fidelity training on the OMat24[72], MPtrj, and sAlex[72] datasets, achieving high accuracy despite heterogeneity in DFT settings, though still focused on inorganic systems. Conversely, both the DPA-3.1-3M and DPA-2.4-7M models implement a multitask training strategy utilizing the OpenLAM dataset collection (refer to Supporting Information LABEL:tab:si-datasets). This collection encompasses a wide array of chemical and material systems, including inorganic materials, catalysis, small molecules, biomolecules, and reactions, all under diverse DFT settings. Detailed hyperparameter information for the models can be found in Supporting Information Table S-1. Unless specified otherwise, the MPtrj task head is employed by default across all benchmarks.
@@ IV.2 Models
GRACE-2L-OAM[74] 12.6M MPtrj No 6.0
IV.3 LAMBench Implementation
@@ IV.4 Generalizability metrics
In this study, we denote the error metric as $M^{m}_{k,p,i}$, where $m$ indicates the model, $k$ denotes the domain index, $p$ signifies the prediction index, and $i$ represents the test set index. For instance, in force field tasks, the domains include Molecules, Inorganic Materials, and Catalysis, such that $k\in\{\text{Molecules, Inorganic Materials, Catalysis}\}$. The prediction types are categorized as energy ($E$), force ($F$), or virial ($V$), with $p\in\{E,F,V\}$. For the specific domain of Molecules, the test sets are indexed as $i\in\{\text{MD22, ANI-1x, AIMD-Chig}\}$. To facilitate a fair comparison, the error metric is truncated and normalized against the error metric of a baseline model (dummy model) as follows:
$\hat{M}^{m}_{k,p,i}=\min\left(\frac{M^{m}_{k,p,i}}{M^{\mathrm{dummy}}_{k,p,i}},1\right)$ (1)
This baseline model predicts energy based solely on the chemical formula, disregarding any structural details, thereby providing a reference point for evaluating the improvement offered by more sophisticated models. For each domain, we compute the log-average of normalized metrics across all datasets within this domain by
@@
For the force field generalizability tasks, we adopt RMSE as the primary error metric. The prediction types include energy and force, with weights assigned as $w_{E}=w_{F}=0.5$. When periodic boundary conditions are assumed and virial labels are available, virial predictions are also considered. In this scenario, the prediction weights are adjusted to $w_{E}=w_{F}=0.45$ and $w_{V}=0.1$. The resulting error metric after averaging over all domains is referred to as $\bar{M}^{m}_{\mathrm{FF}}$.
For the domain-specific property calculation tasks, we adopt the MAE as the primary error metric. In the Inorganic Materials domain, the MDR phonon benchmark predicts maximum phonon frequency, entropy, free energy, and heat capacity at constant volume, while the elasticity benchmark evaluates the shear and bulk moduli. Each prediction type is assigned an equal weight of 1/6. In the Molecules domain, the TorsionNet500 benchmark evaluates the torsion profile energy, torsional barrier height, and the number of molecules for which the predicted torsional barrier height error exceeds 1 kcal/mol. The Wiggle150 benchmark assesses the relative conformer energy profile. Each prediction type in this domain is assigned a weight of 1/4. In the Catalysis domain, the OC20NEB-OOD benchmark evaluates the energy barrier, reaction energy change (delta energy), and the percentage of reactions with predicted energy barrier errors exceeding 0.1 eV for three reaction types: transfer, dissociation, and desorption. Each prediction type in this domain is assigned a weight of 1/5. The resulting error metric after averaging over all domains is denoted as $\bar{M}^{m}_{\mathrm{PC}}$.
IV.5 Applicability metrics
The applicability metrics incorporate both efficiency and stability and are computed differently from the generalizability metrics due to the absence of a dummy model baseline. To evaluate efficiency, we define an efficiency metric, $M_{E}^{m}$ , by normalizing the average inference time (with unit $\mathrm{\mu s/atom}$), $\bar{\eta}^{m}$, of a given LAM measured over 900 configurations with respect to an artificial reference value, thereby rescaling it to a range between zero and positive infinity. A larger value indicates the higher efficiency.
$M_{\mathrm{E}}^{m}=\frac{\eta^{0}}{\bar{\eta}^{m}},\quad\eta^{0}=100\ \mathrm{\mu s/atom},\quad\bar{\eta}^{m}=\frac{1}{900}\sum_{i}^{900}\eta_{i}^{m},$ (5)
@@ IV.5 Applicability metrics
This result is bounded within the range $0,+\infty)$, where a lower value signifies greater stability.
Author contributions
A.P., C.C., L.Z., and H.W. conceived the idea of this work and developed the benchmark workflow. M.G., D.Z., and C.Z. collected and analyzed the results across the benchmark tasks. W.J. and Y.W. performed DFT calculations for dataset relabeling. A.L. provided guidance on the adaptation of the phonon benchmark. C.W. and W.E. provided guidance on the workflow design. All authors contributed to the discussions and revised the manuscript.
@@ Acknowledgments
[…]
@@
Table S-2: Summary of datasets, including the dataset name, number of frames, average number of atoms, data cleaning procedures, DFT level, and dataset weight used in training.
@@
Dataset Name # Frames Avg. # atoms Data Cleaning Level of Theory Weight
@@
Alex2D[77] trn: 1,223,831 val: 135,810 9.9 Exclude frames with energies >> 0 eV/atom, maximum absolute force >> 5 eV/Å, maximum absolute virial/atom >> 8 eV/atom; Trajectory first/last 3 frames + every 10 frame PBE/PAW, 520 eV, 0.4 Å-1 1.83
@@
MPTraj[13] trn: 1,401,956 val: 110,918 31.35 According to the distribution difference between MACE-MP-0 prediction and label, the union set of top 0.1% with the largest uncorrected energy difference and the top 0.1% with the largest force difference were excluded, resulting in a total of 3123 frames excluded. PBE (+U)/PAW, 520 eV, 0.04 Å-1 4.66
OC20M[65] trn: 20,000,000 val: 999,866 73.1 - rPBE/PAW, 350eV 11.65
@@
OC22[78] trn: 8,194,770 val: 394,727 79.8 Exclude frames with energies >> 0eV/atom, maximum absolute force >> 10eV/Å PBE (+U)/PAW, 500 eV 8.14
@@
ODAC23[79] trn: 2,682,332 val: 63,623 203 Exclude frames with energy/atom >> 0.5 eV/atom or << -0.2 eV/atom, and maximum absolute force >> 25 eV/Å; Trajectory first/last 3 frames + every 20th frame PBE-D3/PAW, 600 eV, gamma 1.19
@@
OMAT24[72] trn: 100,568,820 val: 1,074,647 18.6 Exclude frames with energies >> 0 eV/atom, << -25eV/atom, maximum absolute force >> 50eV/Å PBE (+U)/PAW 22.16
[…]
UniPero[82] trn: 14,487 val: 1,357 53.64 - PBEsol/LCAO, 1360 eV, 0.189 Å-1 0.77
Dai2023Alloy[83] trn: 71,482 val: 1,240 20.99 - PBE/Norm-conserving (NC), 1360 eV, 0.094 Å-1 0.67
Zhang2023Cathode[84] trn: 88,692 val: 9,695 46.28 - PBE (+U)/PAW, 520 eV, 0.25 Å-1 1.64
Gong2023Cluster[85] trn: 143,418 val: 15,331 23.67 - PBE-D3/TZV2P, 400-1000 Ry 1.07
Li2025APEX[86] trn: 24,097 val: 100 24.03 - PBE/NC, 1360 eV, 0.15 Å-1 0.42
Shi2024Electrolyte[87] trn: 65,393 val: 3,438 192.69 - PBE-D3, 800 Ry 5.85
Shi2024SSE[88] trn: 125,083 val: 6,587 48.33 - PBE-sol/LCAO, 1360 eV, 0.28 Å-1 2.03
[…]
Li2025General[90] trn: 14,024,587 val: 1,558,260 17.8 Exclude maximum absolute force >> 20eV/Å GFN2-xTB 7.95
Huang2021Deep-PBE[91] trn: 17,582 val: 886 218.59 - PBE/PAW, 650 eV, 0.26 Å-1 1.72
Liu2024Machine[92] trn: 215,481 val: 23,343 70.36 - PBE/LCAO, 1360 eV, 0.151 Å-1 3.88
Zhang2021Phase[93] trn: 46,077 val: 2,342 172.3 - SCAN/PAW, 1500 eV, 0.5 Å-1 4.39
Jiang2021Accurate[94] trn: 138,194 val: 3,965 23.16 - PBE/PAW, 650 eV, 0.1 Å-1 0.51
Chen2023Modeling[95] trn: 6,449 val: 276 16.94 - SCAN/PAW, 650 eV, 0.1 Å-1 0.03
Unke2019PhysNet[96] trn: 2,594,609 val: 136,571 21.38 - revPBE-D3(BJ)/def2-TZVP 4.09
Wen2021Specialising[97] trn: 10,054 val: 474 20.02 - PBE/PAW, 650 eV, 0.1 Å-1 0.05
Wang2022Classical[98] trn: 14,935 val: 738 24.47 - PBE/PAW, 650 eV, 0.1 Å-1 0.07
Wang2022Tungsten[99] trn: 42,297 val: 2,100 24.91 - PBE/PAW, 600 eV, 0.16 Å-1 0.12
Wu2021Deep[100] trn: 27,660 val: 917 96 - PBE/PAW, 600 eV, $\times2$$\times$2 kpt 0.95
Huang2021Deep-PBEsol[91] trn: 7,502 val: 384 160.8 - PBE-sol/PAW, 650 eV, 0.26 Å-1 0.83
Wang2021Generalizable[101] trn: 64,239 val: 2,256 21.85 - PBE-D3/PAW, 650 eV, 0.1 Å-1 0.26
Wu2021Accurate[102] trn: 11,621 val: 568 21.73 - PBE/PAW, 700 eV 0.14
Tuo2023Hybrid[103] trn: 48,078 val: 2,530 45.11 - PBE-D3(BJ)/PAW, 500 eV, 0.16 Å-1 0.59
@@
Table S-3: Phonon related property prediction on the MDR phonon benchmark.
@@
Table S-4: Benchmark results on elasticity related property prediction.
Model MAE GVRH MAE KVRH
DPA-3.1-3M-DomainMatch 9.83 8.68
DPA-3.1-3M 10.77 10.13
DPA-2.4-7M 17.76 16.46
MACE-MP-0 26.20 11.01
MACE-MPA-0 10.27 15.03
Orb-v2 66.07 44.08
Orb-v3 9.75 7.58
SevenNet-l3i5 19.42 9.93
SevenNet-MF-ompa 9.54 9.46
MatterSim-v1-5M 12.75 14.95
GRACE-2L-OAM 9.14 7.46
dummy 67.54 136.26
Table S-5: Torsional MAE and MAEB errors between LAM predictions and its reference DFT labels on the TorsionNet-500 Benchmark.
@@
DPA-3.1-3M-DomainMatch 0.338 0.551 76
DPA-3.1-3M 0.825 1.382 298
DPA-2.4-7M 0.739 1.229 265
MACE-MP-0 1.652 2.437 356
MACE-MPA-0 1.468 2.150 339
Orb-v2 1.248 2.109 346
Orb-v3 0.988 1.421 268
SevenNet-l3i5 1.127 1.641 308
SevenNet-MF-ompa 1.619 2.338 360
MatterSim-v1-5M 1.360 2.234 342
GRACE-2L-OAM 1.394 1.985 309
dummy 2.494 5.818 493
@@
Table S-6: Prediction of conformer relative energies on the Wiggle150 benchmark.
Model MAE (kcal/mol)
DPA-3.1-3M-DomainMatch 1.72
DPA-3.1-3M 5.67
DPA-2.4-7M 14.84
MACE-MP-0 26.60
MACE-MPA-0 14.91
Orb-v2 6.46
Orb-v3 11.92
SevenNet-l3i5 13.88
SevenNet-MF-ompa 10.97
MatterSim-v1-5M 10.73
GRACE-2L-OAM 12.14
dummy 105.06
Table S-7: Energy barrier prediction on the OC20NEB-OOD Benchmark.11footnotemark: 1
Model MAEEa22footnotemark: 2 (eV) MAEdE33footnotemark: 3 (eV) $\phi_{\text{Transfer}}$44footnotemark: 4 $\phi_{\text{Dissociation}}$44footnotemark: 4 $\phi_{\text{Desorption}}$44footnotemark: 4
DPA-3.1-3M-DomainMatch 1.172 0.158 65.1 69.0 59.8
DPA-3.1-3M 1.203 0.234 66.9 63.9 81.1
DPA-2.4-7M 1.271 0.306 68.6 77.9 69.3
MACE-MP-0 1.468 0.580 76.6 84.8 90.6
MACE-MPA-0 1.459 0.566 72.6 84.8 94.5
Orb-v2 2.682 1.729 66.3 74.1 78.7
Orb-v3 2.298 1.470 61.7 72.2 87.4
SevenNet-l3i5 1.423 0.546 77.7 76.0 87.4
SevenNet-MF-ompa 2.070 1.278 66.9 68.4 92.9
MatterSim-v1-5M 2.686 2.018 76.0 84.8 83.5
GRACE-2L-OAM 1.583 0.704 65.1 72.2 90.6
dummy 2.360 0.950 95.4 99.4 95.3
11footnotemark: 1
Only 460 out-of-distribution trajectories were used in this task. 22footnotemark: 2The mean absolute error of the reaction barrier. 33footnotemark: 3The mean absolute error of the reaction energy. 44footnotemark: 4Percentage of reactions with an energy barrier error exceeding 0.1 eV. The analysis considers three reaction types: transfer, dissociation, and desorption.
S-3 Structures for Conservativeness Benchmark
@@
S-4 Inference Efficiency Convergence Test
@@
S-5 The Coordination Number Analysis of Configurations in Efficiency Benchmarks
@@
S-6 Stability Test Results
Table S-8: NVE molecular dynamics simulations over nine atomic systems.